OpenStudy (anonymous):

Rationalize the denominator and simplify. Assume that all variables represent positive real numbers. 3x/(^5sqrt121x^23y^12)

OpenStudy (anonymous):

\[\huge \frac{3x}{\sqrt[5]{121\cdot x^{23} y^{12}}}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whoaa... that's not easy lol

OpenStudy (anonymous):

I know

OpenStudy (anonymous):

example: \[\LARGE \sqrt[5]{3} \cdot x= (\sqrt[5]{3}\;)^5\] do you know what number should be instead of x ?

OpenStudy (anonymous):

\[\LARGE \frac{3x}{\sqrt[5]{121\cdot x^{23}\cdot y^{12}}} \cdot \frac{\left( \sqrt[5]{121\cdot x^{23}\cdot y^{12}}\right)^4 }{\left( \sqrt[5]{121\cdot x^{23}\cdot y^{12}}\right)^4}\] try this ...

OpenStudy (anonymous):

ok